OpenStudy (hellomynameiskam):

What is the solution to the equation? 2/5x= 20 A. 50 B. 8 C. 10 D. 40

OpenStudy (anonymous):

since u need x by itself you need to divide 20 by 2/5

OpenStudy (anonymous):

which equals 50 and that is a

OpenStudy (crystal_bliss):

Simplify 2/5x to 2x/5 2x/5=20 2 Multiply both sides by 5 2x=20×5 3 Simplify 20×5 to 100 2x=100 4 Divide both sides by 2 x=100/2 5 Simplify 100/2 to 50 x=50 50 is your answer

OpenStudy (malikasif):

your qouestion options are wrong

OpenStudy (anonymous):

?????

OpenStudy (crystal_bliss):

no there not wrong because the answer is 50 so it's A

OpenStudy (malikasif):

x =1/50 you can check it by putting x= 1/50 in the equation

OpenStudy (crystal_bliss):

@malikAsif so x= 50 1/50 is same as 50

OpenStudy (anonymous):

no you would check it by multiplying 2/5 by 50 which =50/1 not 1/50

OpenStudy (malikasif):

1/50 = 0.02 not equal to 50

OpenStudy (crystal_bliss):

@HELLOMYNAMEISKAM your answer would be A which is 50.

OpenStudy (malikasif):

how prove your answer

OpenStudy (anonymous):

@Crystal_Bliss yea i agree

OpenStudy (malikasif):

\[\frac{ 2 }{ 5x } = 20 => 2= 5x*20 => 2= 100x => x= \frac{ 2 }{ 100 } => x= \frac{ 1 }{ 50 }\]